OpenStudy (anonymous):

If x² - y² = 24, and x - y = 6, then y =

OpenStudy (alexandervonhumboldt2):

hint: (x-y)*(x+y)=x^2-y^2

OpenStudy (alexandervonhumboldt2):

\[x^2-y^2=(x-y)*(x+y)\]

OpenStudy (poisoned_ruby):

Exactly what I was going to say^^ lol

OpenStudy (alexandervonhumboldt2):

so we have: 6*(x+y)=24 x+y=4 x-y=6 can you solve this system?

OpenStudy (anonymous):

i got -1

OpenStudy (alexandervonhumboldt2):

4-x=x-6 -2x=-2 x=-1 so x is -1 yeah, now substitute -1 for x and solve for y